Jason's bank account contains 30 coins(all nickles or dimes). He has four times as many nickels as dimes. How much money does he have?
Answer by algebrapro18(249)

Well first lets assign varables. Let n be the number of nickles and d be the number of dimes. Then from the problem we get the following system of equations:

n+d=30
4d=n

now since n is already solved for in the second equation then we can just substitute that into the first equation and solve for d.

n+d=30
4d+d=30
d = 6

now we can substitute that back into the second equation and find n.

4d=n
4(6)=n
24=n

so we know that Jason has 24 nickles and 6 dimes for a total of $0.6+$1.20 = $1.80.